# Practitioners’ Perspectives on a Differential Privacy Deployment Registry

### v1

@@ 5.3.2 RQ2b: Informing future deployments
Did not rely on the registry for determining appropriateness of DP. Before making lower-level implementation choices, practitioners must first decide whether DP is appropriate for a given use case. When given a hypothetical use case or considering one of their own (Task 3), participants tended not to rely on the registry for deciding whether DP would be appropriate (P3–5, P7, P11, P12)—despite being explicitly asking them to use the registry for this task. Instead, they relied on their prior knowledge about when DP is well-suited to a task, and considered how well attributes of the use case fit their understanding of when to apply DP. Sometimes, this meant that participants reflected on DP’s appropriateness without returning to the registry. For example, when reading the hypothetical use case, P3 noticed that membership inference attacks [35]—which DP protects against—would be concerning, as knowing someone visited the website described in the use case could be a privacy violation on its own. Based on this observation, they decided DP should be applied for the data release. P4, P7, P11 all noticed that the size of the hypothetical dataset was large enough that DP would be effective (in terms of preserving utility), and similarly made their decision without consulting the registry.
@@
Used “similar” deployments as reference for implementation choices. When asked to talk through how they might use the registry to make implementation choices for either the hypothetical use case described in Task 3 or a use case from their own work, most participants looked to the registry to first find similar deployments (P3, P4, P6, P8, P10–13, P15, P16). However, their understandings of “similarity” varied. For a few participants (P3, P12, P13), similarity was based on the type of data product, like summary statistics or machine learning model. To this end, P3 searched the data product description column for “summary statistics” while P12 searched for “microdata.” For others, similarity was based more on the underlying data to be protected (P8, P10, P11, P13, P15). For instance, as P8 considered the hypothetical use case, they explained that as “the first step, [they] would try to…find deployments in the same domain, which is either surveys with demographic information or relatedly, anything health related.” Participants also focused on the size of the data product (P6, P8), and in P8’s case, both the “scale of the input and also the scale of the output.” Both these participants wished they could filter deployments by dataset size, which the registry does not support.
@@ 5.4 RQ3: Opportunities and challenges
Opportunity: Support broader communication about DP. Participants saw multiple opportunities for using the registry for broader communication about DP. Most notably, they thought the registry could help them advocate within their organizations for the use of DP (P1, P6, P7, P9, P15), particularly to policy or legal teams. For example, P9 stressed that being able to show internal legal teams other examples of deployments would help convince them that DP should be used, given the importance of precedence in legal contexts. P6 and P15 spoke about using the registry in conversations with teammates in managerial roles who presumably do not have deep DP expertise but hold decision-making power. In this way, it seems that participants envision the registry as having potential as a boundary object [36, 37] that can facilitate conversations between people in their organizations with different backgrounds and understandings of DP.
@@ 6 Impact and collaboration
While conducting this research, we worked in collaboration with a larger team at OpenDP, Oblivious, and NIST to create a community-driven deployable DP registry [14]. The work described herein informed the development of the registry intended for deployment; at the same time, there were aspects of mutual influence between that registry and the design described here.
When starting to develop our interface prototype, we used aspects of the prototype created by Oblivious [10] as a starting point. As discussed in the introduction, in spring 2024 a team at Oblivious started to work toward a more systematic, community-driven registry to be further developed in collaboration with OpenDP. They created an initial prototype in blog post format with a table of deployments and educational content about DP. We expanded their basic set of attributes to create a holistic, hierarchical schema. This schema was then incorporated back into their registry, with a few minor modifications. There was mutual influence during the design of the interactive table and deployment card. For example, we shared early versions of an updated prototype and had shared discussions about further iterations, especially with regard to designing the deployment card. Finally, while the content in our “Guide” is largely distinct from the educational content in the initial prototype created by Oblivious, we used their text as a starting point.
Throughout our research process, we shared design ideas, code for the visualizations, our schema, our Tier 2 and Tier 3 entries, and text from our guide with the larger team. Much of our design contributions and filled-out entries have been incorporated back into the registry intended for deployment. This registry is being launched in mid September 2025 alongside a proposal by NIST to host the registry [14].
@@ 7.1 Toward increased adoption of the registry
To maximize the registry’s long-term impact, practitioners must add new deployments as they are made. However, our results suggest that one of the biggest challenges to the registry gaining adoption will be the time and effort required for practitioners to add new entries, especially at the Tier 3 level. To overcome this challenge, we propose strategies to make it easier to add new entries.
@@
First, it may be fruitful to develop partially automated, human-in-the-loop approaches to filling out new entries. In this work, we manually filled out each deployment’s entry, but our preliminary explorations into creating a large language model (LLM) agent for this task appear promising. Such an agent would extract information for the deployment card (from publicly-available documentation, like papers and blog posts), attach supporting quotes for each attribute, and indicate whether the information was explicitly stated, inferred from context, or unspecified, along with page, section, or figure references so that each value can be traced to its source. A practitioner, either involved in the deployment or with adequate DP expertise, would then review and correct the suggestions before submitting to the registry. This review step is essential, as legitimacy of the registry hinges on the correctness of information provided; discrepancies between information published about a deployment outside the registry versus information in the registry could significantly impact the perceived credibility of the registry. Hence, care should be taken to not only verify the agent’s suggestions, but also to improve the agent based on practitioner feedback, for example, by adopting a dual-agent strategy where one model extracts information comprehensively (prioritizing recall) and another extracts information only when highly confident (prioritizing precision), so that overlaps can be accepted with minimal effort while disagreements are flagged for review. Ideally, those maintaining the registry would recommend a single LLM agent for use, and solicit feedback about practitioners’ experiences so that it can be iteratively improved.
@@
Second, it would be valuable for DP libraries and software to automatically save implementation details in a format that can be easily transferred to a deployment card. For example, once a differentially-private program is executed, the practitioner should then be able to download a file with attributes about the deployment (e.g., flavor, definition of adjacent datasets, privacy parameters, composition, mechanisms). Some sociotechnical information, such as the intended use of the data product or rationale for implementation choices, will need to be manually filled out by the practitioner. Extracting information automatically from the implementation code has the added benefit of being able to verify correctness of the claims, which is impossible to do with only a paper describing implementation details.
Third, we draw inspiration from a comment a participant made during our study, and propose making the process of registering a deployment either recommended or mandatory when submitting papers about a real-world DP deployment to peer-reviewed venues, such as Theory and Practice of Differential Privacy (TPDP). Integrating registration into the peer-review process will help build a norm within the community about details that should be shared and increase the chances that a new deployment gets entered into the registry. Other computer science conferences have implemented processes that can serve as guides, such as NeurIPS’ paper checklist [38].
7.2 Fostering community through a discussion board
@@
Participants saw opportunity for the registry becoming a “hub” for the DP community, implying that the registry could become a meeting ground for practitioners. Currently, the registry is designed to be interactive insofar as practitioners can add new deployments to the registry. However, there is significant potential for the registry to facilitate richer forms of interaction between community members. As a basic step, it would be useful for each deployment to list contact information for someone who has agreed in advance to answer questions about the deployment. This simple addition would also help foster an environment where question-asking is encouraged and not limited to only people one already knows. Several participants described how, when working on prior deployments, they often relied on either research papers or consultations (either formally or informally) with experts; for practitioners newer to DP, having contact information available would be especially helpful.
@@
To enable even more discussion, we also propose adding a discussion board to the registry. A discussion board would allow community members to have back-and-forth discussion about deployments. Such a space for discussion would help standards emerge more intentionally. One concern raised by participants is that precedents will drive future decisions, even if these precedents are not advisable. A forum would allow for richer understandings of each deployment, and to what extent its implementation choices should be repeated. In other words, a discussion board would help create more normative guidance or interpretation, which participants in our study seemed to be seeking. Adding a discussion board would, of course, raise a host of moderation challenges. A trial addition of the board can help assess whether its benefits outweigh any potential costs.
@@ 7.3 Making the registry usable beyond practitioners
The original proposal for a registry [1] envisioned it as being useful not only for practitioners, but also for other parties, like policymakers interacting with DP. Our study further suggests that part of the registry’s value is in facilitating communication with policymakers and data subjects. As is, our prototype is designed to support practitioners.
As future work, we envision the registry being expanded to provide tailored support to policymakers, data subjects, and data users in interpreting implementation choices and finding information most relevant to their needs. In particular, the registry can be expanded to include several “modes,” each intended for a different party. For example, the mode for policymakers may provide interpretations of the deployments in their particular legal contexts and make connections between aspects of DP and various laws or policies [39, 40]. Support for data subjects may involve providing lay explanations of the privacy and accuracy guarantees offered by a particular deployment (see [41] for several approaches proposed by prior work). The extent to which these explanations can be automatically generated based on information already in the deployment card requires further research. Finally, when a data product is made publicly-available (i.e., not only published within an organization), it would be helpful to support data users in understanding the implications of added DP noise on their downstream applications. Data curators may submit documentation about their deployments geared toward data users alongside their registry entry. Such documentation may focus specifically on how to account for DP noise in downstream applications [42]. Together, these possible expansions of the registry can increase its impact to a wider range of parties.

### v2

@@ 5.3.2 RQ2b: Informing future deployments
Did not rely on the registry for determining appropriateness of DP. Before making lower-level implementation choices, practitioners must first decide whether DP is appropriate for a given use case. When given a hypothetical use case or considering one of their own (Task 3), participants tended not to rely on the registry for deciding whether DP would be appropriate (P3–5, P7, P11, P12)—despite being explicitly asking them to use the registry for this task. Instead, they relied on their prior knowledge about when DP is well-suited to a task, and considered how well attributes of the use case fit their understanding of when to apply DP. Sometimes, this meant that participants reflected on DP’s appropriateness without returning to the registry. For example, when reading the hypothetical use case, P3 noticed that membership inference attacks shokri2017membership—which DP protects against—would be concerning, as knowing someone visited the website described in the use case could be a privacy violation on its own. Based on this observation, they decided DP should be applied for the data release. P4, P7, P11 all noticed that the size of the hypothetical dataset was large enough that DP would be effective (in terms of preserving utility), and similarly made their decision without consulting the registry.
@@
Used “similar” deployments as reference for implementation choices. When asked to talk through how they might use the registry to make implementation choices for either the hypothetical use case described in Task 3 or a use case from their own work, most participants looked to the registry to first find similar deployments (P3, P4, P6, P8, P10–13, P15, P16). However, their understandings of “similarity” varied. For a few participants (P3, P12, P13), similarity was based on the type of data product, like summary statistics or machine learning model. To this end, P3 searched the data product description column for “summary statistics” while P12 searched for “microdata.” For others, similarity was based more on the underlying data to be protected (P8, P10, P11, P13, P15). For instance, as P8 considered the hypothetical use case, they explained that as “the first step, [they] would try to…find deployments in the same domain, which is either surveys with demographic information or relatedly, anything health related.” Participants also focused on the size of the data product (P6, P8), and in P8’s case, both the “scale of the input and also the scale of the output.” Both these participants wished they could filter deployments by dataset size, which the registry does not support. A future version of the registry may include size of the dataset as an additional attribute of each deployment, ideally measured by approximate number of privacy units.
@@ 5.4 RQ3: Opportunities and challenges
Opportunity: Support broader communication about DP. Participants saw multiple opportunities for using the registry for broader communication about DP. Most notably, they thought the registry could help them advocate within their organizations for the use of DP (P1, P6, P7, P9, P15), particularly to policy or legal teams. For example, P9 stressed that being able to show internal legal teams other examples of deployments would help convince them that DP should be used, given the importance of precedence in legal contexts. P6 and P15 spoke about using the registry in conversations with teammates in managerial roles who presumably do not have deep DP expertise but hold decision-making power. In this way, it seems that participants envision the registry as having potential as a boundary object star1989structure, star1989institutional that can facilitate conversations between people in their organizations with different backgrounds and understandings of DP.
@@ 6 Impact and collaboration
While conducting this research, we worked in collaboration with a larger team at OpenDP, Oblivious, and NIST to create a community-driven deployable DP registry NIST.IR.8588.2025. The work described herein informed the development of the registry intended for deployment; at the same time, there were aspects of mutual influence between that registry and the design described here.
When starting to develop our interface prototype, we used aspects of the prototype created by Oblivious original_registry as a starting point. As discussed in the introduction, in spring 2024 a team at Oblivious started to work toward a more systematic, community-driven registry to be further developed in collaboration with OpenDP. They created an initial prototype in blog post format with a table of deployments and educational content about DP. We expanded their basic set of attributes to create a holistic, hierarchical schema. This schema was then incorporated back into their registry, with a few minor modifications. There was mutual influence during the design of the interactive table and deployment card. For example, we shared early versions of an updated prototype and had shared discussions about further iterations, especially with regard to designing the deployment card. Finally, while the content in our “Guide” is largely distinct from the educational content in the initial prototype created by Oblivious, we used their text as a starting point.
Throughout our research process, we shared our interface, code for the visualizations, our schema, our Tier 2 and Tier 3 entries, and text from our guide with the larger team. Much of our design contributions and filled-out entries have been incorporated back into the registry intended for deployment. This registry was launched in September 2025 with a proposal by NIST to host it NIST.IR.8588.2025.
@@ 7.1 Toward increased adoption of the registry
To maximize the registry’s long-term impact, practitioners must add new deployments as they are made. However, our results suggest that a big challenge to the registry gaining adoption will be the time and effort required for practitioners to add new entries, especially at the Tier 3 level. Thus, we propose strategies to make it easier to add entries.
@@
First, it may be fruitful to develop partially automated, human-in-the-loop approaches for adding new entries. In this work, the first two authors manually filled out each deployment’s entry. We also did some preliminary explorations into creating a large language model (LLM) agent for this task. Such an agent would extract information for the deployment card (from publicly-available documentation, like papers and blog posts), attach supporting quotes for each attribute, and indicate whether the information was explicitly stated, inferred from context, or unspecified, along with page, section, or figure references so that each value can be traced to its source. A practitioner, either involved in the deployment or with sufficient DP expertise, would then review and correct the suggestions before submitting to the registry. Review is essential, as legitimacy and credibility of the registry hinges on information being correct. Care should be taken to not only verify the agent’s suggestions, but also to improve the agent based on practitioner feedback, for example, by adopting a dual-agent strategy where one model extracts information comprehensively (prioritizing recall) and another extracts information only when highly confident (prioritizing precision), so that overlaps can be accepted with minimal effort while disagreements are flagged for review. Ideally, registry maintainers would recommend a single LLM agent, and solicit feedback from practitioners after use so that it can be iteratively improved.
@@
Second, it would be valuable for DP libraries and software to automatically save implementation details in a format that can be easily transferred to a deployment card. For example, once a differentially-private program is executed, the practitioner should then be able to download a file with attributes about the deployment (e.g., flavor, definition of adjacent datasets, privacy-loss parameters, composition, mechanisms). Some sociotechnical information, such as the intended use of the data product or rationale for implementation choices, will need to be manually filled out by the practitioner. Extracting information automatically from the implementation code has the added benefit of being able to verify correctness of the claims, which is impossible to do with only a paper describing implementation details.
Third, based on a participant’s comment, we suggest creating incentives to add new deployments as part of conference publication procedures. For example, if a paper’s contribution is a real-world DP deployment, authors may submit an entry to the registry as part of an optional artifact review, further lending credibility and exposure to their work. This would help incentivize adding entries among academic researchers, and would need to be translated to other types of venues to reach industry practitioners.
7.2 Fostering community through a discussion board.
@@
Participants saw opportunity for the registry becoming a “hub” for the DP community, implying that the registry could become a meeting ground for practitioners. Currently, the registry is designed to be interactive insofar as practitioners can add new deployments to the registry, and users of the registry can interactively explore and compare deployments. However, there is significant potential to facilitate richer forms of interaction between community members. One basic change would be for each deployment to include contact information for someone who has agreed to answer questions about the deployment. This simple addition would also help foster a culture of question-asking, even between people who are not already acquainted. Several participants described how, when working on deployments, they often relied on either research papers or consultations (either formally or informally) with experts; for practitioners newer to DP, contact information would be especially helpful.
@@
To enable further discussion, we propose adding to the registry a discussion board, which would allow community members to have back-and-forth discussions about deployments. These discussions would help standards emerge more intentionally. Participants in our study were concerned that precedents will drive future decisions, even if these precedents are not advisable. A forum would cultivate richer understandings of each deployment, including the extent to which its specific choices should be repeated. Thus, a discussion board would help create more normative guidance or interpretation, which participants in our study seemed to be seeking. Of course, a discussion board would raise a host of moderation challenges, which could be assessed during a trial period to assess benefits and costs.
@@ 7.3 Making the registry usable beyond practitioners
The original proposal for a registry dwork2019differential envisioned it as being useful not only for practitioners, but also for other parties, like policymakers interacting with DP. Our study further suggests that part of the registry’s value is in facilitating communication with policymakers and data subjects. As is, our prototype is designed to support technically skilled practitioners.
As future work, we envision the registry being expanded to provide tailored support to policymakers, data subjects, and data users in interpreting implementation choices. In particular, the registry can be expanded to include several “modes,” each intended for a different party. For example, the mode for policymakers may provide interpretations of the deployments in their particular legal contexts and make connections between aspects of DP and various laws or policies nissim2018privacy, cohen2020towards. Support for data subjects may mean providing lay explanations of the privacy and accuracy guarantees offered by a particular deployment (see dibia2024sok for several approaches proposed by prior work). The extent to which these explanations can be automatically generated based on information already in the deployment card requires further research. Finally, when a data product is made publicly available, it would be helpful to support data users in understanding the implications of added DP noise on their downstream applications. Data curators may submit links to deployment documentation for data users alongside their registry entry. Such documentation may provide guidance on accounting for DP noise in downstream applications triedman2025confidence. Together, these possible expansions can increase the registry’s impact beyond practitioners.
@@ Acknowledgments
Appendix
Pre-task interview guide
  • Please describe your current role.
    • How long have you been in this role?
    • What are your primary responsibilities in this role?
  • How have you interacted with DP either in your current role or prior roles?
  • Have you been involved with making a DP deployment? If yes, then:
    • What kind of data were you seeking to protect? What domain was it in?
    • What were your goals in deploying DP?
    • What was the published data product?
    • What were some challenges you faced?
    • What resources did you use to inform your choices, if any?
    • How did you communicate your deployment with external parties?
  • Have you previously analyzed or evaluated a DP deployment? If yes, then:
    • Describe a DP deployment that you’ve analyzed or evaluated in the past.
    • What were your goals?
    • Which aspects of the deployment did you focus on in your analysis or evaluation?
    • What resources did you use to support your analysis or evaluation, if any?
Post-task interview guide
  • What did you find challenging about using the registry? What did you find easy?
  • What do you see as opportunities the registry could open up?
  • What do you see as challenges to community-wide adoption of the registry?
    • How might a registry gain legitimacy among practitioners?
    • Do you see any potential barriers to new entries being added to the registry? For example, if you were to make a DP deployment in the future, would you add it to the registry?
  • What are some ways you think the value of the registry can be maximized through governance? In other words, how should the registry be governed by the community? How should new deployments be validated, if at all? How should changes to the registry — such as new pieces of information required to add a deployment — be vetted?
  • How do you envision using a registry like this in your work in the future, if at all?
  • Any other thoughts?
Task 3 hypothetical use case
Your organization maintains a website with articles about different medical conditions. Your team has conducted a survey to find out who’s using the website, how they’re using it, and for what purposes.
You have received several tens of thousands of responses. Your team plans to publish aggregate statistics about the collected data.
These published statistics will be used internally at your organization to help improve the website. They will also be made available to researchers interested in studying questions about how people in different demographic or socioeconomic groups access medical information.
Sample survey questions include, but are not limited to:
  • In the last year, how often did you visit this website?
    • More than once a week
    • Once a week
    • Once or twice a week
    • Once a month
    • A few times
    • Did not visit the website in the last year
  • Which purposes do you use the website for? (select all that apply)
    • To find information about medical conditions I have been diagnosed with
    • To find information about medical conditions I have not been diagnosed with
    • To find information about medical conditions friends or family have been diagnosed with
    • To improve my general understanding about various medical conditions
  • Select up to two of the following website sections or features which you use most frequently.
    • Symptom checker
    • Support group finder
    • Diseases & conditions database
    • Drugs & supplements database
    • Tests & procedures database
  • What is your gender? (select all the apply)
    • Female
    • Male
    • Non-Binary
    • Prefer to self describe:
  • Please specify your race / ethnicity. (select all that apply)
    • Hispanic or Latino
    • Black or African American
    • White
    • American Indian or Alaska Native
    • Asian, Native Hawaiian, or Pacific Islander
    • Prefer to self describe:
  • What is the highest level of school you have completed or the highest degree you have received?
    • Less than a high school degree
    • Some high school credit, no diploma or equivalent
    • High school graduate (high school diploma or equivalent including GED)
    • Some college but no degree
    • Associate’s degree
    • Bachelor’s degree
    • Advanced degree (e.g., master’s, doctorate)
  • Which of the following includes your total HOUSEHOLD income for last year, before taxes (in USD)?
    • Less than \$10,000
    • \$10,000 to under \$50,000
    • \$50,000 to under \$100,000
    • \$100,000 to under \$200,000
    • \$200,000 or more
First, use the registry to determine whether DP seems appropriate for this use case.
If you believe DP is appropriate: how might you use the registry to inform decisions about how to implement DP in this use case? What aspects of other deployments would you consider? Which deployments would you focus on the most?
Aspects of the hypothetical website described above were based on WebMD and the Mayo Clinic’s website.